OpenStudy (anonymous):

Find the slope of the straight line that passes through the given points or state that the slope is undefined. Then indicate if the line through the points rises (left to right), falls (left to right), is horizontal, or is vertical 1a. (2, -5) and (-1, -1) 1b. (7, 1) and (7, 2) 1c. (1, 7) and (-2, -2) 1d. (-1, -5) and (-4, -5)

OpenStudy (anonymous):

\[(y2 -y1)/(x2-x1) = m slope \] the slope is undefined if x2-x1 = 0

OpenStudy (anonymous):

Okay, and does it rise or fall ?

OpenStudy (anonymous):

rise over run is how you read the slope fraction, say you have 5/3 you'd read that on the graph as going up 5 units then going to the right or left( depending if the slope is negative or not). After you graph the slope you read a graph from left to right that is how you see if it rises or fall

OpenStudy (anonymous):

Ohhh okay I got it, cool thank you. An what about the other three?

OpenStudy (anonymous):

All you have to do is follow the same steps as long as the question is asking you "Find the slope of the straight line that passes through the given points or state that the slope is undefined. Then indicate if the line through the points rises (left to right), falls (left to right), is horizontal, or is vertical" Its the same steps.

OpenStudy (anonymous):

also, if y2-y1 = 0, and x2-x1 is not 0 you get a horizontal line, if x2-x1 = 0 and y2-y1 is any number, this is division by 0 which gives your slope as undefined which means a vertical line.